Mr. West drove his car from home to Chicago at the rate of 40 mph. and returned at a rate of 45 mph. If his time going exceeded his time returning by 30 minutes, find his time going and his time returning.

time going:
time returning:

Thanks!!

Answer by scott8148(6628)

let x=(time returning), so x+30=(time going) ... d=rt

the distance is the same both ways so 40(x+30)=45x

40x+1200=45x ... 1200=5x ... x=240 (4 hr)

x+30=270 (4.5 hr)